# Review: Quantum Metrology and Sensing with Many-Body Systems

### v1

@@ XIV Global Sensing: Parameter Estimation with Minimal Prior Information
In contrast to local estimation theory, global estimation theory applies when no prior information is available. Therefore, the unknown parameter varies within a broader sensing interval. Consequently, the choice of measurement to be performed on the probe and the analysis of the estimation data should be optimal on average, that is it should be equally effective for any possible value of the parameter Helstrom (1976); Holevo (2011a).
@@
Thermometry, the process of estimating the temperature of a sample, plays a crucial role in quantum thermodynamics Campbell et al. (2017); De Pasquale and Stace (2018) and has various applications, such as determining macroscopic objects in their ground state Chan et al. (2011); Montenegro et al. (2020). In certain scenarios, local estimation for thermometry may not be sufficient, making an alternative estimation strategy highly important. Note that the bound in Eq. (87) may not be saturated as, in general, the optimal measurement setup vary for different values of the unknown parameter $\lambda$. In some cases that the optimal measurement is not a function of $\lambda$ the bound can indeed be saturated. An example of such scenario is thermometry for which energy measurement is optimal, independent of the temperature of the system. This allows to design an optimal probe for measuring temperature over an arbitrary interval. In Ref. Correa et al. (2015), it is shown that optimization of the probe reduces to optimizing the energy eigenvalues. Interestingly, for the case of local thermometry, where temperature varies over a narrow interval, the optimal probe contains only two distinct eigenenergies, with a unique ground state and full degeneracy for the rest of the spectrum. The energy gap between the ground state and the rest of the spectrum depends on the temperature Paris (2015)and can be obtained by solving a nonlinear equation Correa et al. (2015). Using the global sensing formulation of Eq. (87), one can show that the optimal probe requires more distinct energy eigenvalues as the temperature varies over a larger interval Mok et al. (2021). In fact, through exploiting an evolutionary algorithm the authors of Ref. Mok et al. (2021) find out that an effective two-level probe for local sensing turns into a three-level system with the two lowest energy eigenvalues being unique and then the rest of the spectrum gets fully degenerate as the temperature interval increases. By further enhancing the interval more and more energy levels are separated from the degenerate band. It is worth emphasizing that the bounds, given by Eq. (87) can be saturated for all these optimal thermometers.
@@
While both the van Trees and the average uncertainty approaches focus on minimizing the variance of estimation, a new formalism for global thermometry has been developed based on minimization of the mean logarithmic error. By exploiting scale invariance and other well-behaved properties, such as symmetric invariance and monotonicity increment (decrement) from (towards) the actual hypothesis temperature value Rubio et al. (2021), the mean logarithmic error $\bar{\epsilon}_{\mathrm{mle}}$ is defined as the global thermometry figure of merit when the temperature varies over a prior distribution
@@
XV Summary and Outlook
@@ 2. Optimal measurement parameter-value
Despite current progress on many-body quantum sensing, there are still several open problems to explore. While criticality has been identified as a resource for sensing in equilibrium scenarios, the criteria which results in quantum enhanced sensing is not specified in non-equilibrium probes. Another problem which requires further investigation is the performance of quantum sensors under imperfect situations, such as the presence of decoherence. In general terms, while quantum sensors are naturally expected to be sensitive to the desired parameters, we would like them to be robust against other system’s parameters or noisy imperfections. Despite several works on noisy quantum metrology Escher et al. (2011b); Kołodyński and Demkowicz-Dobrzański (2013); Rossi et al. (2020); Demkowicz-Dobrzański and Maccone (2014); Demkowicz-Dobrzański et al. (2017); Dür et al. (2014); Nichols et al. (2016); Chaves et al. (2013); Alushi et al. (2024); Faist et al. (2023); Reiter et al. (2017), the notion of robustness has not yet been formulated quantitatively for quantum sensors. Furthermore, a general issue for quantum sensors arise in the multi-parameter Cramér-Rao inequality as the bounds are not tight and thus saturating them may not be achievable. Developing tighter bounds and strategies towards achieving them in many-body sensors require closer connections between quantum metrology and control theory. One important problem in multi-parameter sensing is the situation where Fisher information matrix is singular and thus sensing becomes impossible. Specifying the criteria which results in singularity of the Fisher information matrix in many-body systems and recipes to fix them requires further attention. Finally, experimental realization of quantum many-body sensors is a big objective which ultimately defines the success of these probes. Many physical platforms can implement many-body sensors, including nuclear magnetic resonance systems, Rydberg atoms, optical lattices, superconducting simulators, nitrogen vacancies and levitated objects. progress on these setups as well as new physical platforms open further opportunities for developing many-body quantum sensors with wide range of applications.
@@ Acknowledgements.
AB acknowledges support from the National Natural Science Foundation of China (Grants No. 12050410253, No. 92065115, and No. 12274059), and the Ministry of Science and Technology of China (Grant No. QNJ2021167001L). VM acknowledges funding by the National Natural Science Foundation of China (Grants No. 12050410251 and No. 12374482). RY acknowledges support from the National Natural Science Foundation of China for the International Young Scientists Fund (Grant No. 12250410242).

### v2

@@ XIV Global Sensing: Parameter Estimation with Minimal Prior Information
In contrast to local estimation theory, global estimation theory applies when no (or some) prior information is available. Recently, the impact of uncertainty in control parameters on the sensitivity of critical sensors has been explored Mihailescu et al. (2024c). Therefore, in these situations the unknown parameter varies within a broader sensing interval. Consequently, the choice of measurement to be performed on the probe and the analysis of the estimation data should be optimal on average, that is it should be equally effective for any possible value of the parameter Helstrom (1976); Holevo (2011a).
@@
Thermometry, the process of estimating the temperature of a sample, plays a crucial role in quantum thermodynamics Campbell et al. (2017); De Pasquale and Stace (2018); Zhang and Tong (2022) and has various applications, such as determining macroscopic objects in their ground state Chan et al. (2011); Montenegro et al. (2020). In the context of previously discussed equilibrium quantum sensing, it has been demonstrated that finite temperature can surprisingly enhance the performance of critical quantum metrology protocols Ostermann and Gietka (2024). In addition, in certain scenarios, local estimation for thermometry may not be sufficient, making an alternative estimation strategy highly important. Note that the bound in Eq. (87) may not be saturated as, in general, the optimal measurement setup vary for different values of the unknown parameter $\lambda$. In some cases that the optimal measurement is not a function of $\lambda$ the bound can indeed be saturated. An example of such scenario is thermometry for which energy measurement is optimal, independent of the temperature of the system. This allows to design an optimal probe for measuring temperature over an arbitrary interval. In Ref. Correa et al. (2015), it is shown that optimization of the probe reduces to optimizing the energy eigenvalues. In Ref. Campbell et al. (2018), this analysis was extended to investigate the spectral features required for sensitivity to multiple temperatures. Interestingly, for the case of local thermometry, where temperature varies over a narrow interval, the optimal probe contains only two distinct eigenenergies, with a unique ground state and full degeneracy for the rest of the spectrum. The energy gap between the ground state and the rest of the spectrum depends on the temperature Paris (2015)and can be obtained by solving a nonlinear equation Correa et al. (2015). Using the global sensing formulation of Eq. (87), one can show that the optimal probe requires more distinct energy eigenvalues as the temperature varies over a larger interval Mok et al. (2021). In fact, through exploiting an evolutionary algorithm the authors of Ref. Mok et al. (2021) find out that an effective two-level probe for local sensing turns into a three-level system with the two lowest energy eigenvalues being unique and then the rest of the spectrum gets fully degenerate as the temperature interval increases. By further enhancing the interval more and more energy levels are separated from the degenerate band. It is worth emphasizing that the bounds, given by Eq. (87) can be saturated for all these optimal thermometers.
@@
While both the van Trees and the average uncertainty approaches focus on minimizing the variance of estimation, a new formalism for global thermometry has been developed based on minimization of the mean logarithmic error. By exploiting scale invariance and other well-behaved properties, such as symmetric invariance and monotonicity increment (decrement) from (towards) the actual hypothesis temperature value Rubio et al. (2021), the mean logarithmic error $\bar{\epsilon}_{\mathrm{mle}}$ is defined as the global thermometry figure of merit when the prior information about the temperature is wide
@@
XV Optimal Control and Machine Learning Strategies for Quantum Sensing
Let us recall that the ultimate precision attainable through any quantum sensing algorithm depends on four steps (see Fig. 2): (i) initialization of the probe, (ii) evolution of the probe under the influence of the system’s (known and unknown) parameters, subjected to external controls, (iii) extraction of information from the probe’s state through a suitably chosen measurement, and (iv) estimation of the unknown parameter value from the extracted measurement statistics. We have so far reviewed results where these steps were all assumed to be perfectly achievable. However, for practical sensing tasks, three valuable resources are inevitably limited: (i) limited probe size available for sensing, (ii) finite coherence timescale of the probe available for initialization and evolution, and (iii) finite measurement time leading to degradation of measurement accuracy and precision with each round. Attempts to lift these restrictions have been partially discussed in Secs. XI.1-XI.2-XIII. In general, performing a constrained optimization over all these valuable yet limited resources is analytically intractable and numerically demanding in all but the simplest cases. So far, broadly two approaches have been successfully employed in this direction, which we briefly review here.
XV.1 Optimal Control Based Approaches
Quantum optimal control theory (QOCT) is now an area of extensive literature. We refer the interested reader to the review in Ref. Koch and Budich (2022) for a more detailed treatment of QOCT in other areas of quantum technology. For sensing purposes within QOCT approach, an extra control Hamiltonian is added to the probe’s evolution. If the probe evolves through a Hamiltonian $H_{\lambda}$ containing information about the unknown parameter $\lambda$, then the time-dependent total Hamiltonian is given by
$H_{tot}(t)=H_{\lambda}+H_{c}(\textbf{u}(t),t),$ (93)
where the control Hamiltonian $H_{c}$ is a function of control parameters $\textbf{u}(t)$. The goal is to design the control Hamiltonian $H_{c}(\textbf{u}(t),t)$ in such a way that maximum precision in estimation of the parameter $\lambda$ is achieved. This can be broadly accomplished in two ways: (quasi)analytically and numerically. In terms of control strategies, broadly two distinct strategies are possible, viz., open-loop control and closed-loop, i.e., feedback control.
XV.1.1 Open-loop QOCT strategies
The simplest and oldest systematic approach to QOCT is to apply Pontryagin’s Maximum Principle (PMP), which builds on the calculus of variations to derive a set of necessary conditions for optimal control, see Ref. Boscain et al. (2021) for a detailed introduction to PMP in quantum settings. PMP introduces an auxiliary system described in terms of a costate $|\mu(t)\rangle$ which must satisfy the original evolution equation backwards in time, i.e., $i\hbar\frac{|d\mu(t)\rangle}{dt}=-H_{tot}(t)|\mu(t)\rangle$ with the same boundary conditions. Now if the task is to optimize a cost-function $\mathcal{C}$, which can be generally written as a sum of two cost functions, i.e., $\mathcal{C}=\mathcal{C}_{1}+\mathcal{C}_{2}$, with $\mathcal{C}_{1}$ being a cost function in terms of some suitable path-dependent quantity integrated over the trajectory of evolution (say, the average QFI for sensing tasks) and $\mathcal{C}_{2}$ being a terminal cost function (say, the fidelity achieved with a target state etc); then PMP puts the necessary condition for the control parameters being optimized, i.e., $\textbf{u}(t)=\textbf{u}^{\text{opt}}(t)$, as the optimization over the following function for every time $t$ and the quantum state $|\psi(t)\rangle$ along the original evolution equation
$\textbf{u}^{\text{opt}}(t)=\arg\max_{\textbf{u}(t)}\left[\langle\mu(t)|H_{tot}|\psi(t)\rangle-\mathcal{C}\right].$ (94)
For very simple problems, this optimization is sometimes possible to perform analytically. In Ref. Lin et al. (2021), the authors consider such a problem, where the twist and turn angular momentum Hamiltonian $H=\chi J_{z}^{2}+\omega J_{z}$ with unknown parameter $\omega$ is augmented with a control Hamiltonian $\Omega(t)J_{x}$ and find the optimal trajectory of $\Omega(t)$ by maximizing the QFI of the final state as the cost function. However, for more complicated problems like those involving many-body interactions, analytically solving Eq. (94) is no longer presumed to be viable. Thus numerical algorithms must be adopted. Popular gradient-based optimization algorithms for quantum control include gradient ascent pulse engineering (GRAPE) Khaneja et al. (2005) and Krotov’s method Reich et al. (2012). In the former, updates are made simultaneously for all times $t$, while in the latter updates are made sequentially. However, for exponentially growing Hilbert space dimensions in generic many-body systems, gradient-based algorithms are highly limited. Moreover, in noisy environments gradient updates are highly unreliable in the first place. Thus gradient-free optimization algorithms are often required. Primary approach along this direction is to Fourier-decompose the control pulse sequence into a smaller number of bases. Randomly sampling a small number of bases via chopped random basis (CRAB) Caneva et al. (2011) thus avoids the dimensionality problem for many-body systems.
XV.1.2 Closed-loop QOCT strategies
The second approach is the so called feedback-based strategies, where measurement results obtained in each round are fed into the quantum control for the next round. Pang and Jordan Pang and Jordan (2017) considered a phase estimation problem for a parameter $\lambda$ imprinted into a time-dependent Hamiltonian $H(\lambda,t)$. The goal is to optimize the QFI along the entire trajectory. To do so, they noted that QFI at each instant $T$ is upper bounded by the following integral
[…]
where $\nu_{\max}(t)$ and $\nu_{min}(t)$ are respectively maximum and minimum eigenvalues of $\partial_{\lambda}H(\lambda,t)$ at any time $t$. Thus, Pang and Jordan employed an additional control Hamiltonian $H_{c}(t)$ to steer the trajectory towards the eigenspace of $\partial_{\lambda}H(\lambda,t)$ as much as possible. The crucial aspect is that the control Hamiltonian is independent of the actual parameter $\lambda$, but written in terms of $\tilde{\lambda}$, a known prior estimate of $\lambda$ which is updated at each round. As an illustration, they considered a qubit in a uniformly rotating magnetic field $\textbf{B}(t)=B(\cos\omega t~{}\hat{x}+\sin\omega t~{}\hat{y})$, where the goal is to estimate $B$ or the field rotation frequency $\omega$. For estimating $\omega$ from evolution up to time $T$, their approach furnishes a QFI $I_{Q}(\omega)=B^{2}T^{4}$. In contrast, in the absence of the control Hamiltonian, QFI scales only quadratically with time. This result indicates that adaptive control strategies are eminently useful for quantum sensing.
Yang et al in Ref. Yang et al. (2020) addressed a slightly different question, preparation of the optimal probe for maximum precision via adaptive feedback control. They used purity loss instead of QFI for ease of direct measurement. Their approach circumvents the problem of exponentially growing Hilbert spaces by avoiding any computation involving matrix representation of the time-independent Hamiltonian $H$. In particular use for many-body sensing, they considered a spin-chain probe with the control being limited to a sequence of local rotations and even the rotation strength can be adjusted only a fixed number of times. Their results, experimentally implemented on an NMR platform, indicate that Heisenberg scaling is possible even with such a severe limitation on the control.
XV.2 Machine Learning Based Approaches
The ongoing revolution in machine learning has already had a significant impact in the study of many-body quantum physics Carleo and Troyer (2017); Gray et al. (2018); Vicentini (2021). These range from identification of various exotic phases of matter in hitherto analytically challenging models Carrasquilla and Melko (2017); Greplova et al. (2020), to Hamiltonian learning Wang et al. (2017), and investigation of foundational issues like Bell-nonlocality Deng (2018). Machine-learning algorithms, which learn from previously acquired data, also assist in solving optimization problems that arise in quantum sensing. For a general overview of machine-learning assisted quantum metrology, we refer the interested reader to the recent review by Huang et al Huang et al. (2024). As we describe below, quantum sensing has already been shown as an especially fruitful area of application for machine learning methods.
XV.2.1 Classical machine learning strategies
Critical state preparation is of practical relevance for quantum sensing in many-body systems, where the enhanced sensitivity of critical states to external fields is leveraged to perform more precise measurements. Reinforcement learning (RL) methods have started showing significant promise in this task, since it does not assume any prior knowledge and only has a cumulative reward function which it seeks to optimize. RL was first employed to optimize over driving protocols to suitably initialize a nonintegrable quantum many-body probe in minimum time and high fidelity by Bukov et al in Ref. Bukov et al. (2018). This was specialised for matrix product states (MPS) by Metz and Bukov in Ref. Metz and Bukov (2023), which shows that RL can adapt to control protocols in real time in the presence of stochastic perturbations as well. Meanwhile, RL has also identified nonlinear measurements as another resource for state preparation in optical cavities Porotti et al. (2022). For the estimation part, RL was used for general time-dependent sensing problems by Xu et al Xu et al. (2019) where the reward function is the QFI given by the control, and by Xiao et al Xiao et al. (2022), where they employ a reward function in terms of the ultimate QFI bound as well as design a general reward function. Control pulse engineering via a RL algorithm, viz., the Asynchronous Advantage Actor-Critic (A3C) algorithm, has been demonstrated for atomic spin ensembles Qiu et al. (2022), resulting in attainment of Heisenberg limit. In yet another approach, Cimini et al Cimini et al. (2023) employed RL to train on the experimental data to learn Bayesian updates in a multiparameter setting. These general approaches show RL is competitive with GRAPE or other optimal control methods while being more resource-efficient, robust, and not being restricted by veracity of prior assumptions. Thus RL in particular holds much promise for being adapted specifically to many-body sensing probes in future.
Apart from RL, one can exploit machine learning as an estimation algorithm. In Ref. Nolan et al. (2021a), the authors explore regression for inferring a machine-learning estimation of an unknown parameter. The method is necessarily frequentist, i.e. relying on repeated estimates to build up statistics, and eventually converges to the maximum likelihood Bayesian estimation algorithm. In a similar approach, parameter estimation can be considered as a sequence of classification tasks, for which one can use artificial neural networks to efficiently perform Bayesian estimation Nolan et al. (2021b).
XV.2.2 Quantum machine learning strategies
The above approaches seek to solve the quantum metrological control problem through classical machine-learning algorithms. Thus, a natural question arises: can we use recent advances in quantum machine-learning algorithms for improved sensing? Variational quantum algorithms, seeking to optimize a cost function via parametric quantum circuits, is perhaps the most important of these advances. Variational quantum metrology (VQM) is a direct offshoot of these algorithms. The generic recipe for VQM in many-body systems is to build up the Hamiltonian metrological evolution as a sequence of Trotter steps represented by the quantum circuits and then consider a QFI-like cost function to optimize over. Koczor et al. Koczor et al. (2020) used VQM to search for the optimal probe state for an interferometry task. Similar works were reported in Refs. Ma et al. (2021); MacLellan et al. (2024) and carried over to the multiparameter setting (Kaubruegger et al., 2023; Le et al., 2023), with a complete treatment available in Ref. Meyer et al. (2021) and experimental photonic implementation in Ref. Cimini et al. (2024). Experimentally, programmable quantum sensors based on VQM has been demonstrated for ion-trap Marciniak et al. (2022), superconducting qubit Gong et al. (2023), and Rydberg interaction based neutral atomic array Kaubruegger et al. (2019) platforms, with concrete tasks like building atomic clocks Kaubruegger et al. (2021) promising to become very important. Specialized computational packages like MetQuan are also becoming available Sinha et al. (2024) recently.
XVI Summary and Outlook
@@ 2. Optimal measurement parameter-value
Despite current progress on many-body quantum sensing, there are still several open problems to explore. While criticality has been identified as a resource for sensing in equilibrium scenarios, the criteria which results in quantum enhanced sensing is not specified in non-equilibrium probes. Another problem which requires further investigation is the performance of quantum sensors under imperfect situations, such as the presence of decoherence. In general terms, while quantum sensors are naturally expected to be sensitive to the desired parameters, we would like them to be robust against other system’s parameters or noisy imperfections. Despite several works on noisy quantum metrology Escher et al. (2011b); Kołodyński and Demkowicz-Dobrzański (2013); Rossi et al. (2020); Demkowicz-Dobrzański and Maccone (2014); Demkowicz-Dobrzański et al. (2017); Dür et al. (2014); Nichols et al. (2016); Chaves et al. (2013); Alushi et al. (2024); Faist et al. (2023); Reiter et al. (2017); Mihailescu et al. (2023, 2024a), the notion of robustness has not yet been formulated quantitatively for quantum sensors. Furthermore, a general issue for quantum sensors arise in the multi-parameter Cramér-Rao inequality as the bounds are not tight and thus saturating them may not be achievable. Developing tighter bounds and strategies towards achieving them in many-body sensors require closer connections between quantum metrology and control theory. One important problem in multi-parameter sensing is the situation where Fisher information matrix is singular and thus sensing becomes impossible. Specifying the criteria which results in singularity of the Fisher information matrix in many-body systems and recipes to fix them requires further attention. Finally, experimental realization of quantum many-body sensors is a big objective which ultimately defines the success of these probes. Many physical platforms can implement many-body sensors, including nuclear magnetic resonance systems, Rydberg atoms, optical lattices, superconducting simulators, nitrogen vacancies and levitated objects. progress on these setups as well as new physical platforms open further opportunities for developing many-body quantum sensors with wide range of applications.
@@ Acknowledgements.
AB acknowledges support from the National Natural Science Foundation of China (Grants No. 12050410253, No. 92065115, and No. 12274059), and the Ministry of Science and Technology of China (Grant No. QNJ2021167001L). VM acknowledges funding by the National Natural Science Foundation of China (Grants No. W2432005, No. 12374482, and No. 12050410251). RY acknowledges support from the National Natural Science Foundation of China for the International Young Scientists Fund (Grant No. 12250410242).